OpenStudy (anonymous):

[1 , infinity) except 1 How would you represent this new domain into simpler interval form?

OpenStudy (anonymous):

\[1\le x <\infty \]

OpenStudy (anonymous):

but interval notation is what u have already written

OpenStudy (anonymous):

Except" 1 "

OpenStudy (anonymous):

F(x) = 2 + (sqrt x -1) G(x) = sqrt (x -1)

OpenStudy (anonymous):

Find the domain of f/g

OpenStudy (anonymous):

ok .........then put an open interval at one \[(1,\infty)\]

OpenStudy (anonymous):

But I kind get confused because of that denominator which theorem will I follow ?

OpenStudy (anonymous):

n/ Sqrt (x-1) > Which theorem will I follow is it the not equal to zero or greater than zero

OpenStudy (anonymous):

he domain of a function is the set of "input" or argument values for which the function is defined. That is, the function provides an "output" or value for each member of the domain.[1] The set of values the function may take is termed the range of the function.

OpenStudy (anonymous):

for \[\sqrt{x-1}\] the domain is \[x-1\ge 0\] but since in this case sqrt(x-1) is in denominator,so it cannot be equal to zero.ie.domain is x-1>0 x>1 \[(1,\infty)\] @Reyjuanx10

OpenStudy (anonymous):

So if it is in the denominator automaticilly it must be not equal to zero even if it has square root?

OpenStudy (anonymous):

Can I apply rationalizing ?

OpenStudy (anonymous):

ya .........i ll give an example:: say the function is 1/(x-2) now domain is the possible set of values x can have......... now x cannot be equal to 2 as if x=2 the value of function will be infinite which is not possible.... so domain in this case will be all real numbers except 2 ... square implies that the nos.....have to be positive....right

OpenStudy (anonymous):

this was the reply of ur denominator question

OpenStudy (anonymous):

Got it . Thanks !

OpenStudy (anonymous):

ur welcome.....